# Refinable modeling for unbinned SMEFT analyses

### v1

@@ 7.5 An analytic toy example
${\textrm{d}}\sigma(\boldsymbol{x}|\nu_{1},\nu_{2})$ $=N\exp\left(0.25\left(\nu_{1}\sin(x)+\nu_{2}\cos(0.5x)\right)^{2}\right){\textrm{d}}x$ (107)
with support $x\in[-\pi,\pi]$. The logarithm of the cross-section is a quadratic polynomial in ${\boldsymbol{\nu}}$ for all $\boldsymbol{x}$, suggesting a perfect fit with a two-parameter parametric tree at quadratic accuracy. For the training, we chose five base points $\mathcal{V}=\{(0.5,0),\,(0,0.5),\,(1,0),\,(0,1),\,(0.5,0.5)\}$ that lead to a full-rank matrix $V$ in Eq. 104. With a nominal data set at $(\nu_{1},\nu_{2})=(0,0)$, we have six synthetic data sets, each with $5\cdot 10^{5}$ events, sufficient to train the algorithm. We fit $B=100$ boosting iterations and require a maximum tree depth of $4$ and a minimum requirement for the number of events in each terminal node, which is 50 events. The learning rate is set to 0.2 for all boosting iterations.
Figure 3: The DCR for various model parameters for the toy study described in the text. Dashed lines show the true DCR from the training data; solid lines show the prediction obtained from the parametric regression tree algorithm. The top (bottom) panel compares after $B=5$ ($B=100$) boosting iterations. The bottom panel also includes new model parameter points not present during training. In this panel, the prediction is drawn after the truth, and the dashed histograms are not separately visible because of the good quality of the fit with $B=100$.
@@ 8 Top quark pair production in the $2\ell$ channel
As an example process, we now study dileptonic top quark pair production in proton-proton collisions at $\sqrt{s}=13~{}\,\text{TeV}$, $\textrm{pp}\rightarrow\mathrm{t}\overline{\mathrm{t}}\rightarrow\textrm{b}\ell^{+}\nu_{\ell}\overline{\textrm{b}}\ell^{-}\overline{\nu}_{\ell}$. At the parton and particle level the event simulators provide access to all the needed quantities. For the calibration of reconstructed objects (jets, missing transverse momentum, and leptons), the ATLAS and CMS open data projects [63, 64] make available information on uncertainties in the calibration. It is sufficient to showcase the application of the tools. However, a fully realistic detector simulation with a comprehensive assessment of all systematic effects, dependent on the data-taking period, is neither possible nor needed for the following demonstrator. We restrict the study to a heuristic treatment of the main uncertainties of the differential cross-section measurement. A binned measurement of $\mathrm{t}\overline{\mathrm{t}}(2\ell)$, including a comprehensive account of systematic uncertainties, is available from the ATLAS [65] and CMS [66] Collaborations.
@@ 8.1 Event simulation
Using the Delphes objects, we reconstruct the top quark kinematic quantities described in Ref. [66]. This provides access to SMEFT-sensitive observables, including the top quarks’ invariant masses, angles, and transverse momenta. To reduce the computational demand while keeping sensitivity to SMEFT effects, we require $m(\mathrm{t}\overline{\mathrm{t}})>750$  GeV, corresponding to an inclusive fiducial cross-section of 0.31 pb [66]. We normalize the Delphes simulation of a total of $1.2\times 10^{6}$ events to this value and use a central value for the luminosity $\mathcal{L}_{0}=137\,$fb-1 with a conservative 5% log-normal uncertainty,
@@
We denote the nominal sample by $\mathcal{D}_{0}$ and the phase-space region after the event selection by $\mathcal{X}$.
@@
where the left and right-chiral lower-case quark fields $q$ and $u$ belong to the first and second generation. The third-generation left-chiral quark doublet is denoted by $Q$. A non-zero value of $C_{tG}$ provides CP-even and CP-odd modifications to the top-gluon interaction. The four fermion operators add contact interactions of the first and second with the third-generation quark currents. For obtaining the SMEFT predictions, we use the reweighting technique discussed in Sec. 4.4.3 and Ref. [7].
@@
The following event-level features define the observation $\boldsymbol{x}$. From the reconstructed top quark momenta, we compute the invariant mass $m(\mathrm{t}\overline{\mathrm{t}})$, the transverse momentum $p_{\textrm{T}}({\textrm{$\mathrm{t}\overline{\mathrm{t}}$}})$, the rapidity difference $\Delta\eta(\mathrm{t}\overline{\mathrm{t}})=\eta(\textrm{t})-\eta(\overline{\textrm{t}})$ and the difference of absolute rapidities of the top and anti-top quark, $\Delta|\eta|(\mathrm{t}\overline{\mathrm{t}})=|\eta(\textrm{t})|-|\eta(\overline{\textrm{t}})|$. The quantities $m(\mathrm{t}\overline{\mathrm{t}})$ and $p_{\textrm{T}}({\textrm{$\mathrm{t}\overline{\mathrm{t}}$}})$ are sensitive to SMEFT effects with energy-growth while $\Delta|\eta|(\mathrm{t}\overline{\mathrm{t}})$ is sensitive to the effects of the charge-asymmetry [75, 76], modified, e.g., by $C_{tj}^{(8)}$. Furthermore, we include the transverse momentum and the pseudo-rapidity of the top and anti-top quark. Because leptons are clean probes of the possible SMEFT effects, independent of the hadronic activity, we also include the invariant mass $m(\ell^{+}\ell^{-})$, the transverse momentum $p_{\textrm{T}}(\ell^{+}\ell^{-})$, the rapidity difference $\Delta\eta(\ell^{+}\ell^{-})$, and the difference of absolute rapidities $\Delta|\eta|(\ell^{+}\ell^{-})$ of the dilepton system. Finally, we include the absolute value of the difference of the azimuthal lepton angles $|\Delta\varphi_{\textrm{lab}}|(\ell^{+}\ell^{-})$ and the cosine of the spatial angle between the leptons $\cos(\phi_{\textrm{lab}}(\ell^{+}\ell^{-}))$ as measured in the lab frame. The distributions of these observables for the SM and non-zero values for the Wilson coefficients are shown in Fig. 4. We find good agreement with the study in Ref. [8]. The CMS measurement of the spin-correlation in $\mathrm{t}\overline{\mathrm{t}}$ [77] found constraining power for $C_{tG}$ in the distribution of products of angular observables of the leptons, measured in a specific reference frame spanned by the top quark momentum and the beam plane [78]. These variables characterize the spin-density matrix of the $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ system, and we present a brief description and their distributions in Appendix C. The resulting distributions for non-zero values of the Wilson coefficients are shown in Fig. 4.
@@
The estimate of the detector-level SMEFT dependence of the signal process $\hat{R}(\boldsymbol{x}|{{\boldsymbol{\theta}}})$ can be learned by one of the tools in Refs. [8, 9, 10, 11, 12, 13, 15, 16, 17, 14]. We use the Boosted Information Tree technique [9, 10] to learn the polynomial SMEFT dependence up to the quadratic order. Concretely, we train trees with a maximum depth of four in $B=300$ boosting iterations with a learning rate of $\eta=0.2$. We regularize each tree by requiring at least 50 events in each node.
@@
We use the parametric tree from Sec. 7 to estimate the systematic effects discussed in the following sections. Similar settings turn out to be almost universally applicable. We keep the maximum tree depth at four in all cases and use $B=300$ boosting iterations and a learning rate of $\eta=0.2$. When we obtain the synthetic data from reweighting, such that there are no independent stochastic fluctuations in the various terms in the loss function, a minimum node size requirement of 50 events proves sufficient to regularize the trees. For systematic variations where $\boldsymbol{x}$ changes with ${\boldsymbol{\nu}}$, we raise this regulator requirement to 500.

### v3

@@ 5.4 The Boosted Parametric Tree algorithm
$\hat{\Delta}_{A,j}^{(b)}=\left[\;\sum_{\nu\in\mathcal{V}}{\boldsymbol{\nu}}{\boldsymbol{\nu}}^{\textrm{T}}\right]_{AB}^{-1}\left[\sum_{\;{\boldsymbol{\nu}}\in\mathcal{V}}{\boldsymbol{\nu}}\log\frac{\sigma^{(b)}_{j,{\boldsymbol{\nu}}}}{\sigma^{(b)}_{j,{\mathbf{0}}}}\right]_{B}.$ (61)
The inverse matrix in the first factor exists if $\mathcal{V}$ has a full-rank coordinate matrix, so we need at least as many training samples ${\boldsymbol{\nu}}\in\mathcal{V}$ as there are coefficient functions $\hat{\Delta}_{A}(\boldsymbol{x})$. This linear relation of log-ratios makes Eq. 61 highly efficient to evaluate. We can then use the CART or “Tree Alternate Optimization” (TAO) [55, 56, 57, 58] algorithms to determine the optimal phase-space partitioning $\mathcal{J}^{(b)}$, completing iteration $b$. After $B$ boosting iterations, all constants in Eq. 57 are determined, giving the final DCR estimate
$\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})=\exp(\nu_{A}\hat{\Delta}_{A}(\boldsymbol{x}))\simeq\frac{{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})\;\;\;}{{\textrm{d}}\sigma(\boldsymbol{x}|\textrm{SM})}$ (62)
It is fast to evaluate, parametric in ${\boldsymbol{\nu}}$, satisfies $\hat{S}(\boldsymbol{x}|{\mathbf{0}})=1$, and is continuous in both $\boldsymbol{x}$ and ${\boldsymbol{\nu}}$. Algorithm 1 provides a pseudo-code summary of the steps, defining the BPT algorithm. Appendix C contains a detailed derivation and a simple analytic toy example.
6 Gradually refinable modeling
With the setup for the training of parametric regressors in place, we discuss gradually refinable modeling as a flexible approach to unbinned analyses, incorporating incremental improvements while preserving existing results. Based on procedures commonly employed in binned analyses, we discuss the factorization of POI and nuisance parameter dependencies in the unbinned case. Uncorrelated groups of systematic effect are isolated and subsequently learned independently. With the help of an additive model, summing over the various contributing processes, we can incrementally extend and refine an existing model without invalidating existing components. In this way, gradually refinable modeling aligns with established practices for systematic uncertainty management in binned analyses, extending these strategies to unbinned data while allowing the analysis to evolve with growing data and improved modeling techniques.
6.1 The binned Poisson likelihood
We begin with the binned Poisson likelihood for several observations ($N_{\textrm{bin}}$) in disjoint phase-space regions (bins), a setup described in detail in Refs. [59, 60]. Multiple processes, labeled by $p=1,\ldots,N_{p}$, contribute to the cross-section component $\sigma_{n,p}(\textrm{SM})$ in bin $n$. The dependence on SMEFT POIs is assumed to factorize from the systematic effects. Since SMEFT ME-squared terms are polynomial or can be truncated to polynomial form, a small set of non-zero values of ${{\boldsymbol{\theta}}}$ suffices to determine the coefficients in the SMEFT parametrization $\sigma_{n,p}(\theta)=R_{n,p}({{\boldsymbol{\theta}}})\sigma_{n,p}(\textrm{SM})$ [7]. Here, $R_{n,p}({{\boldsymbol{\theta}}})$ satisfies $R_{n,p}(\textrm{SM})=1$ and fully encodes the SMEFT dependence in each bin. If $R_{n,p}({{\boldsymbol{\theta}}})=1$ holds to a good approximation for all $n$, we call the process $p$ a background.
The Poisson expectation of the yield in bin $n$ can then be expressed as
The binned ansatz in Eq. 63 reflects inductive bias. First, systematic effects are modeled in a factorized form, meaning the constants $\Delta_{n,p,1/2}$ are assumed to be accurately determined by individually varied simulations, with all other model parameters held fixed. Second, the model is additive. This inconspicuous fact, combined with the factorization of systematic effects, is key to enabling gradual model refinement, an implicit feature in the binned case. Once the per-process expectations and systematic parametrizations in Eq. 63 are established, most of these values can remain unchanged even if the model is refined to include a new process or nuisance parameter. Although this computational saving is modest in the binned case, the unbinned analysis replaces the bin-by-bin constants in Eq. 63 with $\boldsymbol{x}$-dependent ML parametrizations. Selecting a flexible additive model minimizes the need for re-training when refining the model, offering potentially significant gains in computational efficiency.
The additivity in Eq. 63 naturally accommodates per-process nuisances related to normalization uncertainties with two key applications. First, higher-order perturbative corrections (“k-factors”), derived from theoretical predictions, enhance the accuracy of inclusive parton-level predictions. These $k$-factors generally apply to a single process, with reduced uncertainties best captured by nuisances that scale only this component. Second, normalization nuisances are useful for small backgrounds where the pdf in $\mathcal{D}$ can be estimated from $\mathcal{A}$, but normalization uncertainties remain significant. In this case, a normalization nuisance allows for in-situ constraints from $\mathcal{D}$. Specifically, setting $\Delta_{n,p,1}=\log\alpha_{\textrm{norm},\,p}$ and $\Delta_{n,p,2}=0$ for all $n$ results in a scaling of process $p$, where $\alpha_{\textrm{norm},\,p}$ is a positive constant that normalizes the impact of the nuisance $\nu_{\textrm{norm},\,p}$. Additionally, we can omit $\nu_{\textrm{norm},\,p}$ from the penalty, allowing the process’s normalization to float during profiling.
6.2 Approximate factorization of systematic effects
We now substitute the DCR in Eq. 6 with an ML surrogate model. “Likelihood-free” inference refers to techniques that rely on parametrically evaluating ratios of the extended likelihood, and thus ratios of the differential cross-section $\textrm{d}\Sigma({{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$. These alone are enough to evaluate Eq. 10.
Constructing a generic ML surrogate starts by expressing the unbinned model $\textrm{d}\Sigma({{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ as a sum over weighted sub-processes, with normalization uncertainties treated separately [^5] The reason for the separate treatment of normalization nuisances is best seen in comparing the Taylor expansion in ${\boldsymbol{\nu}}$ with the corresponding expansion of the purely multiplicative model in Ref. [17] where nuisances are modeled relative to the total differential cross-section instead of per-process. For arbitrary values of normalization nuisances, a polynomial expansion of the logarithm of the total differential cross-section requires arbitrarily many terms that would have to be learned individually. The ansatz in Eq. 64 will reduce the ensuing ML task to a straightforward classification problem, one for each process.. Nuisance parameters $\nu_{p,\textrm{norm}}$ are introduced for this purpose. We have
where event samples for each component ${\textrm{d}}\sigma_{p}(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ can be obtained from Eq. 23, Eq. 30, or Eq. 31. Next, we factorize systematic effects and POI dependence. The SM point is at ${\boldsymbol{\theta}}$ =${\boldsymbol{\nu}}$ =$\mathbf{0}$, and for each ${\textrm{d}}\sigma_{p}(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ we have
This factorization works if SMEFT effects are independent of systematic effects, i.e.,
The factor
approximates SMEFT variations as a polynomial in ${\boldsymbol{\theta}}$ and can be obtained from methods in Refs. [8, 9, 10, 11, 12, 13, 15, 16, 17, 14]. Systematic effects are parametrized by
We can learn this parametric dependence, one effect at a time, using the strategy in Sec. 5. The ML model can be a neural network or the tree-based algorithm from Sec. 5.4.
The validity of Eq. 66 must be established on a case-by-case basis and can be verified through simulation. The separation of particle and detector-level effects from SMEFT effects is generally accurate due to the different energy scales involved; the POIs typically do not influence low-energy detector interactions. At the parton level, we need to verify the independence of POIs from systematic effects. PDFs, for example, may depend on SMEFT POIs [61], so this correlation should not be neglected without careful consideration. Similarly, the linear and quadratic SMEFT terms may have scale uncertainties differing from the SM prediction. In this case, $\hat{S}_{p}(\boldsymbol{x}|\nu_{R},\nu_{F})$ should be trained with synthetic scale variations that cover scale variations for non-zero POIs. A suitably flexible model should accommodate these subtle analysis-dependent effects, which we leave to future treatment. From now on, we assume the factorization
holds accurately. Following the same steps, we factorize $\hat{S}_{p}(\boldsymbol{x}|{\boldsymbol{\nu}})$ into uncorrelated groups of systematic uncertainties and train each factor with Eq. 53. For instance, uncorrelated one-parameter systematic uncertainties with quadratic accuracy simplify the surrogate to
with $2K$ real-valued functions $\hat{\Delta}_{p,k,1}(\boldsymbol{x})$ and $\hat{\Delta}_{p,k,2}(\boldsymbol{x})$ for each $p$. In most cases, first or second-degree polynomials are sufficient, though the method allows higher degrees.
6.3 A general unbinned surrogate model
In analogy to Eq. 63, we define a general model for the fiducial differential cross-section,
Next, we create a likelihood-free ML surrogate, relying only on differential cross-section ratios. Dividing by
gives per-process DCRs that we replace with surrogates. The simplest approach divides each ${\textrm{d}}\sigma_{p}(\boldsymbol{x}|\textrm{SM})$ in Eq. 71 by Eq. 72,
$\frac{{\textrm{d}}\Sigma(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})}{{\textrm{d}}\Sigma(\boldsymbol{x}|\textrm{SM})\;}$ $=\sum_{p=1}^{N_{p}}\,\hat{R}_{p}(\boldsymbol{x}|{{\boldsymbol{\theta}}})\,\alpha_{\textrm{norm},\,p}^{\nu_{\textrm{norm},\,p}}\,\hat{S}_{p}(\boldsymbol{x}|{\boldsymbol{\nu}})\,\hat{g}_{p}(\boldsymbol{x})\quad\textrm{where}\quad\hat{g}_{p}(\boldsymbol{x})\simeq\frac{{\textrm{d}}\sigma_{p}(\boldsymbol{x}|\textrm{SM})}{\sum_{q}{\textrm{d}}\sigma_{q}(\boldsymbol{x}|\textrm{SM})}$ (73)
estimates the DCR of each process relative to the SM total.
A classifier trained to distinguish process $p$ from the total SM simulation can learn $\hat{g}_{p}(\boldsymbol{x})$ using the likelihood ratio trick. The DCR for arbitrary denominators, as required for profiling in Eq. 1, can then be obtained from double ratios, allowing ${\textrm{d}}\Sigma_{p}(\boldsymbol{x}|\textrm{SM})$ to cancel out.
Equation 73 provides significant modeling flexibility, as the quotient of Eq. 71 and Eq. 72 can be represented in various ways using the surrogates $\hat{g}_{p}(\boldsymbol{x})$. We present two examples demonstrating how this flexibility can solve common challenges in analysis development.
6.4 Refining an existing model
New systematic effects can be gradually incorporated, as expanding the dimension of the nuisance vector ${\boldsymbol{\nu}}$ in, for example, Eq. 70 does not invalidate existing surrogates $\hat{\Delta}_{p,1/2}$. Only the new components require training.
Similarly, additional background sources can be seamlessly included. With the additive structure of Eq. 71, a new process can be added by adjusting the steps leading to Eq. 73. For instance, if a missing background ${\textrm{d}}\sigma_{\textrm{BKG}}(\boldsymbol{x})$, with $R_{p}(\boldsymbol{x}|{{\boldsymbol{\theta}}})=1$, is identified, we can extend ${\textrm{d}}\Sigma(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ by adding a term,
We can express the DCR in terms of the existing model as
where
The only new component is $\hat{g}^{\prime}(\boldsymbol{x})$, a classifier that gives the DCR for the new process relative to the previous total. The new process adds to both the numerator and denominator, with no change to the rest of the model. If the new background has uncertainties, we replace ${\textrm{d}}\sigma_{\textrm{BKG}}(\boldsymbol{x})$ with ${\textrm{d}}\sigma_{\textrm{BKG}}(\boldsymbol{x}|{\boldsymbol{\nu}})$ in Eq. 74 and repeat the derivation, yielding
where $g^{\prime}(\boldsymbol{x})$ from Eq. 76 remains, and we only need to learn one extra factor, $\hat{S}_{\textrm{BKG}}(\boldsymbol{x}|{\boldsymbol{\nu}})$, to model the background’s systematic effects,
similar to Eq. 68. Refinable modeling thus avoids retraining existing regressors and enables incremental analysis development. Because an event sample for ${\textrm{d}}\sigma_{\textrm{BKG}}$ is the only ingredient for obtaining $\hat{g}^{\prime}(\boldsymbol{x})$, it could alternatively be measured in real-data side bands, supporting the development of data-driven unbinned estimation strategies.
6.5 Refinement for a high-purity process
Now, consider a single SMEFT-dependent signal process ${\textrm{d}}\sigma_{\textrm{SMEFT}}(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ and multiple ${\boldsymbol{\theta}}$-independent backgrounds, ${\textrm{d}}\sigma_{p}(\boldsymbol{x}|{\boldsymbol{\nu}})$. To form the DCR, we divide each term in the sums of both sides of Eq. 73 by ${\textrm{d}}\sigma_{\textrm{SMEFT}}(\boldsymbol{x}|\textrm{SM})$, which leaves the result unchanged but modifies the parametrization to
$\frac{{\textrm{d}}\Sigma(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})}{{\textrm{d}}\Sigma(\boldsymbol{x}|\textrm{SM})\;}$ $\simeq\frac{\alpha_{\textrm{norm}}^{\nu_{\textrm{norm}}}\;\hat{R}(\boldsymbol{x}|{{\boldsymbol{\theta}}})\;\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})+\sum_{p=1}^{N_{p}}\,\alpha_{\textrm{norm},\,p}^{\nu_{\textrm{norm},\,p}}\,\hat{S}_{p}(\boldsymbol{x}|{\boldsymbol{\nu}})\,\hat{g}^{\prime\prime}_{p}(\boldsymbol{x})}{1+\sum_{p}\hat{g}^{\prime\prime}_{p}(\boldsymbol{x})}$ (79)
where
and signal quantities have no process index. The classifier $\hat{g}^{\prime\prime}(\boldsymbol{x})$ is trained to distinguish each background from the SMEFT signal at the SM point. Adding a new background process only requires training one additional classifier, as expanding the sum over $p$ leaves the rest of the model unaffected.
The steps in this and the previous section can be combined and repeated as modeling is refined. Equation 71 thus supports incremental refinements, similar to the binned model in Eq. 63. Once a new effect prediction is available, it can be incorporated. Unlike in the binned case, expectations here come from separately trained surrogates rather than binned yields.
@@ 7 Top quark pair production in the $2\ell$ channel
As an example, we study dileptonic top quark pair production in pp collisions at $\sqrt{s}=13~{}\,\text{TeV}$, $\textrm{pp}\rightarrow\mathrm{t}\overline{\mathrm{t}}\rightarrow\textrm{b}\ell^{+}\nu_{\ell}\overline{\textrm{b}}\ell^{-}\overline{\nu}_{\ell}$, or $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ for short. The event simulators provide all necessary quantities at the parton and particle levels. For calibration of reconstructed objects (jets, missing transverse momentum, and leptons), ATLAS and CMS open data projects [62, 63] supply uncertainty information. This is enough to demonstrate the tools’ application. A fully realistic detector simulation with all data-dependent systematic effects is neither feasible nor needed. We focus on a heuristic treatment of the main uncertainties in the differential cross-section. A detailed binned measurement of $\mathrm{t}\overline{\mathrm{t}}(2\ell)$, including a full account of systematic uncertainties, is available from ATLAS [64] and CMS [65]. We focus on a heuristic treatment of the main uncertainties in the differential cross-section. A detailed binned measurement of $\mathrm{t}\overline{\mathrm{t}}(2\ell)$, including a full account of systematic uncertainties, is available from ATLAS [64] and CMS [65].
@@ 7.1 Event simulation
Using the Delphes objects, we reconstruct the top quark kinematic quantities described in Ref. [65]. This provides access to SMEFT-sensitive observables, including the top quarks’ invariant masses, angles, and transverse momenta. To reduce the computational demand while keeping sensitivity to SMEFT effects, we require $m(\mathrm{t}\overline{\mathrm{t}})>750$  GeV, corresponding to an inclusive fiducial cross-section of 0.31 pb [65]. We normalize the Delphes simulation of a total of $1.2\times 10^{6}$ events to this value and use a central value for the integrated luminosity $\mathcal{L}_{0}=137\,$fb-1 with a conservative 5% log-normal uncertainty,
@@
where the left and right-chiral lower-case quark fields $q$ and $u$ belong to the first and second generation. The third-generation left-chiral quark doublet is denoted by $Q$. A non-zero value of $C_{tG}$ provides CP-even and CP-odd modifications to the top-gluon interaction. The four fermion operators add contact interactions of the first and second with the third-generation quark currents. For obtaining the SMEFT predictions, we use the reweighting technique discussed in Sec. 4.4.3 and Ref. [7]. The dominant effect of these operators on the $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ cross-section is linear with the Wilson coefficients [74] so that complications from dominantly quadratic predictions that violate Wilk’s theorem when computing the distribution of the profile likelihood test statistic can be neglected [27].
@@
The following event-level features define the observation $\boldsymbol{x}$. From the reconstructed top quark momenta, we compute the invariant mass $m(\mathrm{t}\overline{\mathrm{t}})$, the transverse momentum $p_{\textrm{T}}({\textrm{$\mathrm{t}\overline{\mathrm{t}}$}})$, the rapidity difference $\Delta\eta(\mathrm{t}\overline{\mathrm{t}})=\eta(\textrm{t})-\eta(\overline{\textrm{t}})$ and the difference of absolute rapidities of the top and anti-top quark, $\Delta|\eta|(\mathrm{t}\overline{\mathrm{t}})=|\eta(\textrm{t})|-|\eta(\overline{\textrm{t}})|$. The quantities $m(\mathrm{t}\overline{\mathrm{t}})$ and $p_{\textrm{T}}({\textrm{$\mathrm{t}\overline{\mathrm{t}}$}})$ are sensitive to SMEFT effects with energy-growth while $\Delta|\eta|(\mathrm{t}\overline{\mathrm{t}})$ is sensitive to the effects of the charge-asymmetry [75, 76], modified, e.g., by $C_{tj}^{(8)}$. Furthermore, we include the transverse momentum and the pseudo-rapidity of the top and anti-top quark. Because leptons are clean probes of the possible SMEFT effects, independent of the hadronic activity, we also include the invariant mass $m(\ell^{+}\ell^{-})$, the transverse momentum $p_{\textrm{T}}(\ell^{+}\ell^{-})$, the rapidity difference $\Delta\eta(\ell^{+}\ell^{-})$, and the difference of absolute rapidities $\Delta|\eta|(\ell^{+}\ell^{-})$ of the dilepton system. Finally, we include the absolute value of the difference of the azimuthal lepton angles $|\Delta\varphi_{\textrm{lab}}|(\ell^{+}\ell^{-})$ and the cosine of the spatial angle between the leptons $\cos(\phi_{\textrm{lab}}(\ell^{+}\ell^{-}))$ as measured in the lab frame. The distributions of these observables for the SM and non-zero values for the Wilson coefficients are shown in Fig. 4. We find good agreement with the study in Ref. [8]. The CMS measurement of the spin-correlation in $\mathrm{t}\overline{\mathrm{t}}$ [77] found constraining power for $C_{tG}$ in the distribution of products of angular observables of the leptons, measured in a specific reference frame spanned by the top quark momentum and the beam plane [78]. These variables characterize the spin-density matrix of the $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ system, and we present a brief description and their distributions in Appendix D. The resulting distributions for non-zero values of the Wilson coefficients are shown in Fig. 4.
@@
The estimate of the detector-level SMEFT dependence of the signal process $\hat{R}(\boldsymbol{x}|{{\boldsymbol{\theta}}})$ can be learned by one of the tools in Refs. [8, 9, 10, 11, 12, 13, 15, 16, 17, 14]. We use the Boosted Information Tree technique [9, 10] to learn the polynomial dependence up to the quadratic order. Concretely, we train trees with a maximum depth of four in $B=300$ boosting iterations with a learning rate of $\eta=0.2$. We regularize each tree by requiring at least 50 events in each node.
@@
We use the parametric tree from Sec. 5.4 to estimate the systematic effects discussed in the following sections. Similar settings turn out to be almost universally applicable. We keep the maximum tree depth at four in all cases and use $B=300$ boosting iterations and a learning rate of $\eta=0.2$. When we obtain the synthetic data from reweighting, such that there are no independent stochastic fluctuations in the various terms in the loss function, a minimum node size requirement of 50 events proves sufficient to regularize the trees. For systematic variations where $\boldsymbol{x}$ changes with ${\boldsymbol{\nu}}$, we raise this regulator requirement to 500.